# Discrete Approximate Circle Bundles

### v2

@@ 6 Coordinatization
Suppose $\mathcal{U}=\{U_{j}\}_{j=1}^{n}$ is a finite open cover of a metric space $B$ and $\Omega\in\check{Z}_{\varepsilon}^{1}(\mathcal{U};\underline{O(2)})$ witnesses that $\{\varphi_{j}\}_{j=1}^{n}\in T_{\alpha}(\pi,\mathcal{U})$ for some map $\pi:X\to B$ ($X$ a finite metric space). Let $\mathcal{W}=\{W^{r}\}$ denote the associated weights filtration of $\mathcal{N}(\mathcal{U})$. For each $r$, we show how to construct a coordinatization map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ for $d\leq 2n$ which is in some sense compatible with $\Omega|_{W^{r}}$. This gives a menu of options for models for the dataset. In the special case when $\Omega|_{W^{r}}$ is an approximate coboundary, we construct a ’global trivialization’ map $F^{r}:X\to B\times\mathbb{S}^{1}$.
@@
    Step 1: (Choose A Filtration Stage) Choose a filtration stage $W^{r}\in\mathcal{W}$ of $\mathcal{N}(\mathcal{U})$. Use Construction 5.5 to construct the sets $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j=1}^{n}$, then choose a partition of unity $\{\rho^{r}_{j}\}_{j=1}^{n}$ subordinate to $\mathcal{U}^{r}$.
@@
    Step 2: (Dimensionality Reduction) For each $j$, define $\widehat{\Phi}^{r}_{j}:U^{r}_{j}\to V(2,2n)$ by
    $\widehat{\Phi}^{r}_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho^{r}_{0}(b)}\Omega_{0j}\\ \vdots\\ \sqrt{\rho^{r}_{n}(b)}\Omega_{nj}\\ \end{array}\right)$ (23)
    Choose a value $d\leq 2n$ and apply Principal Stiefel Coordinates to the point cloud $X_{V}=\bigcup_{j}\Phi_{j}^{r}(U_{j}^{r})\subset V(2,2n)$ to obtain a map $\text{psc}:X_{V}\to V(2,d)$.
@@
    Step 3: (Classifying Map) For each $j$, let $\widetilde{\Phi}_{j}^{r}=\text{psc}\circ\widehat{\Phi}_{j}^{r}:U_{j}^{r}\to V(2,d)$. Define $\tilde{f}^{r}:B^{r}\to\mathbb{R}^{d\times d}$ by
    $\widetilde{f}^{r}(b)=\sum_{j=1}^{n}\rho^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)\widetilde{\Phi}^{r}_{j}(b)^{T}$ (24)
    and let $f^{r}=\Pi\circ\tilde{f}^{r}:B^{r}\to Gr(2,d)$.
@@
    Step 4: (Cocycle Projection) For each $j$, define $\Phi^{r}_{j}:U^{r}_{j}\to V(2,d)$ by $\Phi^{r}_{j}(b)=\widehat{\Pi}(f^{r}(b),\widetilde{\Phi}^{r}_{j}(b))$, where $\widehat{\Pi}$ is the map from Construction 3.15. Then, define $\Omega^{r}\in\check{Z}^{1}(\mathcal{U}^{r};\mathcal{C}_{O(2)})$ by $\Omega^{r}_{jk}(b)=(\Phi^{r}_{j}(b))^{T}\Phi^{r}_{k}(b)$.
@@
    $F_{j}^{r}(x)=\Phi^{r}_{j}(\pi(x))\exp\left(\sum_{k=0}^{n}\rho_{k}^{r}(\pi(x))\log\Omega_{jk}^{r}(\pi(x))\varphi_{k}(x)\right)$ (25)
@@ 6.1 Coordinatization Pipeline For Trivial Bundles
In the special case when a family $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U}^{r})$ is witnessed by a discrete approximate coboundary $\Omega$, we construct a global trivialization map $\varphi:X\to B\times\mathbb{S}^{1}$. In particular, suppose $\omega=\text{det}_{*}(\Omega)\in\check{Z}^{1}(\mathcal{U}^{r};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}=\Delta_{\omega}(\Omega)\in\check{Z}^{2}(\mathcal{U}^{r};\mathbb{Z}_{\omega})$ are coboundaries.
@@
    Construct a solution to the linear system $\delta^{0}\phi=\omega$ over $\mathbb{Z}_{2}$. Define $\hat{\Omega}=\phi\cdot\Omega\in\check{Z}^{1}(\mathcal{U}^{r};\underline{SO(2)})$, and let $\hat{f}_{j}=\phi_{j}\circ f_{j}$ for all $j$.
@@
Since trivial bundles are flat, we can often use Singer’s method [WS13] in practice to obtain a locally-constant true coboundary $\widetilde{\Omega}$ which is close to an approximate one. Though we have no theoretical guarantees on stability, the fact that the result is locally-constant guarantees that the resulting trivialization map does not have unnecessary variation.
7 Implementations
Here we demonstrate the practical viability and utility of the algorithms we have described in this paper through a series of implementations on real and synthetic datasets. In each case, we used the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations. A summary of the results of each experiment can be found in Table 2. Note that the witness quality $\alpha$ in each case was well outside the range required to meet our theoretical hypotheses, which suggests that our approach is stable for a larger class of objects than the current results apply to. Lastly, further details about our implementation on optical flow data can be found in an upcoming companion paper.
@@ 7 Implementations
Prism Densities $\mathbb{RP}^{2}$ 15,408 25,000 $\neq 0$ $\pm 3$
Star Densities $\mathbb{S}^{2}$ 15,408 25,000 0 $\pm 10$
Table 1: A summary of the setups for the three experiments.
Experiment $\varepsilon_{\text{trivs}}$ $\delta$ $\alpha$ $\varepsilon_{\text{cocycle}}$ $d_{Z}(\Omega,\Pi_{Z}(\Omega))$
Optical Flow Patches 0.976 0 0.976 0 0
Prism Densities 0.582 0.272 0.799 0.250 0.248
Star Densities 1.120 0.620 2.949 0.501 0.471
Table 2: Summary of experimental results.
@@
Optical Flow Patches 12 0.976 0.976 0.976 0.976 0.976 12/12 0/0
Prism Densities 73 0.987 0.987 0.231 0.987 0.231 70/216 4/73
Star Densities 146 1.939 1.939 1.939 1.939 1.912 431/432 286/288
Table 3: Summary of persistence information for the computed characteristic class representatives in each experiment.
@@ 7.1 Optical Flow Patches
Optical flow refers to the apparent motion of objects between consecutive frames of a video [HS81]. This perceived motion is characterized by an assignment of a vector $v\in\mathbb{R}^{2}$ to each pixel in each frame. The vector at pixel $(x,y)$ in frame $t$ points to the location in frame $t+1$ where pixel $(x,y)$ appears to move. The estimation of optical flow from a video is central to many computer vision tasks, including motion tracking, object segmentation, and video compression [MG15], [Zha+21]. The Sintel dataset [But+12], derived from the open-source animated short film “Sintel” by the Blender Foundation, was created as a high-quality benchmark for testing and training optical flow models.
@@
In [Ada+20], Adams et al. studied the nonlinear statistics of 3 x 3 high-contrast optical flow patches from [But+12]. They proposed that a dense core subset $X$ of (normalized) high-contrast patches is concentrated around a 2-manifold in $\mathbb{R}^{18}$ homeomorphic to $\mathbb{T}^{2}$. To support their claim, they introduced a feature map $\pi_{\text{flow}}:X\to\mathbb{RP}^{1}$ which roughly measures a patch’s predominant axis of flow (see Figure 4). They then offered evidence that the data in $X$ over any small interval in $\mathbb{RP}^{1}$ has circular topology, clearly supporting the hypothesis that an underlying circle bundle structure is present.
@@
By Proposition 2.13, there are exactly two non-isomorphic circle bundles over $\mathbb{RP}^{1}\cong\mathbb{S}^{1}$ which can be distinguished computationally by their Stiefel-Whitney classes in $H^{1}(\mathbb{RP}^{1};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$. The total spaces of the two bundles are the torus and the Klein bottle. To confirm the torus model proposed by [Ada+20], we used the same pre-processing to obtain a sample of 50,000 normalized high-contrast optical flow patches from the Sintel dataset and refine to a sample $X_{\text{flow}}$ of 25,000 patches from the dense core subset $X(300,50)$. We then computed the predominant axis of flow $\pi_{\text{flow}}(x)\in\mathbb{RP}^{1}$ for each patch $x\in X_{\text{flow}}$. Figure 5 shows the persistence diagram for a random subsample of 1000 points from $X_{\text{flow}}$. For comparison, we also computed persistence diagrams for a random 1000-point sampling from $X(300,30)$ and a noisy synthetic sampling of the proposed torus model. Observe that the synthetic dataset exhibits the Betti signature of $\mathbb{T}^{2}$ ($\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$) over a large range of scales, while the most persistent Betti signature for the samples of $X(300,50)$ and $X(300,30)$ is $\beta_{0}=1$, $\beta_{1}=1$, $\beta_{2}=0$.
Figure 5:
On the other hand, as observed in [Ada+20], the local picture of $X_{\text{flow}}$ with respect to $\pi_{\text{flow}}$ tells a different story. We covered $\mathbb{RP}^{1}$ with 12 open intervals of the form $U_{j}=(\frac{j\pi}{12}-\frac{\pi}{16},\frac{j\pi}{12}+\frac{\pi}{16})$ (note that $\frac{\pi}{16}$ is the maximal radius for which the nerve of $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ has no triple intersections). Figure 6 shows the first two principal components and the persistence diagrams for the data in each $\pi_{\text{flow}}^{-1}(U_{j})$, strongly suggesting the presence of an underlying circle bundle structure.
@@
We computed a family $\{f_{j}\}_{j=1}^{12}$ of discrete approximate local trivializations for $X_{\text{flow}}$ subordinate to $\mathcal{U}$ using the sparse circular coordinates algorithm, then obtained a witness $\Omega\in\check{C}^{1}(\mathcal{N}(\mathcal{U});O(2))$ using the procedure outlined in Construction 3.5. Figure 7 shows a visualization of $\mathcal{N}(\mathcal{U})$; the weight $w(\Omega_{jk})$ of each edge $(jk)$ is shown in black, and the values assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)$ are shown in red. A 0-cochain $\phi\in C^{0}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ satisfying $\delta^{0}\phi=\omega$ is shown in blue, confirming a topological torus structure underlying the data.
Figure 7: (Left) Correlations on intersections between the discrete approximate local trivializations $\{f_{j}\}_{j=1}^{12}$ computed for $X_{\text{flow}}$. (Right) The nerve of the open cover $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ of $\mathbb{RP}^{1}$. Each edge is labeled by its associated weight. The components of the Stiefel-Whitney class representative $\omega=\text{det}_{*}(\Omega)$ are shown in blue, and a potential $\phi$ for $\omega$ is shown in red.
Once we verified the global triviality of the underlying bundle structure, we used the coordinatization pipeline described in Section 6 to obtain a global coordinate system for $X_{\text{flow}}$. To our surprise, we discovered that a significant portion of the patches had ambiguous predominant flow directions despite being high-contrast (see Figure 8). Such patches are not accounted for in the torus model proposed by [Ada+20]. Upon further investigation, we found that the fibers of the predominant direction map $\pi$ are geometrically more like cylinders than circles, and that this extra degree of freedom locally parametrizes variations in directional strength among patches in $X_{\text{flow}}$. We present the details of our unexpected findings and a refined model for the dataset in an upcoming paper. The generalized model we describe also accounts for other families of high-contrast patches not featured in the original torus model which are particularly relevant for motion tracking analysis.
@@
Figure 9: (Left) The optical flow torus model proposed by [Ada+20]. (Right) A sample of coordinatized patches from $X_{\text{flow}}$ with high directional strength, arranged by predominant flow direction (x) and assigned fiber coordinate (y).
From a modeling perspective, the results of our analysis of $X_{\text{flow}}$ demonstrate how the interplay between local and global measurements depends crucially on the geometry of the chosen feature map. On the other hand, a careful fiberwise analysis helped us to both confirm an existing model and identify subtle, informative additional structure in the data which had previously been missed. See our upcoming paper for additional details.

### v3

@@ 6 Coordinatization
Suppose $\mathcal{U}=\{U_{j}\}_{j=1}^{n}$ is a good finite open cover of a metric space $B$ and $\Omega\in\check{Z}_{\varepsilon}^{1}\left(\mathcal{U};\underline{O(2)}\right)$ witnesses that $\{f_{j}\}_{j=1}^{n}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some function $\pi:X\to B$ ($X$ a finite metric space). Let $\mathcal{W}^{\Omega}=\{W^{r}\}$ denote the associated weights filtration of $\mathcal{N}(\mathcal{U})$. For each $r$, we show how to construct a coordinatization map $F^{r}:X\to V(2,d)\times_{O(2)}\mathbb{S}^{1}$ for $d\leq 2n$ which is in some sense compatible with $\Omega|_{W^{r}}$. This gives a menu of options of models for the dataset. In the special case when $\Omega|_{W^{r}}$ is an approximate coboundary, we construct a ’global trivialization’ map $F^{r}:X\to B^{r}\times\mathbb{S}^{1}$.
@@
    Step 1: (Choose A Filtration Stage) Choose a filtration stage $W^{r}\in\mathcal{W}^{\Omega}$ of $\mathcal{N}(\mathcal{U})$. Use Construction 5.5 to construct the sets $\mathcal{U}^{r}=\{U_{j}^{r}\}_{j=1}^{n}$, then choose a partition of unity $\{\rho^{r}_{j}\}_{j=1}^{n}$ subordinate to $\mathcal{U}^{r}$.
@@
    Step 2: (Dimensionality Reduction) For each $j$, define $\widehat{\Psi}^{r}_{j}:U^{r}_{j}\to V(2,2n)$ by
    $\widehat{\Psi}^{r}_{j}(b)=\left(\begin{array}[]{ccc}\sqrt{\rho^{r}_{0}(b)}\Omega_{0j}\\ \vdots\\ \sqrt{\rho^{r}_{n}(b)}\Omega_{nj}\\ \end{array}\right)$ (31)
    Choose a value $d\leq 2n$ and apply Principal Stiefel Coordinates [14] to the point cloud $X_{V}=\bigcup_{j}\Phi_{j}^{r}(U_{j}^{r})\subset V(2,2n)$ to obtain a map $\text{psc}:X_{V}\to V(2,d)$.
@@
    Step 3: (Classifying Map) For each $j$, let $\widetilde{\Psi}_{j}^{r}=\text{psc}\circ\widehat{\Psi}_{j}^{r}:U_{j}^{r}\to V(2,d)$. Define $\widetilde{g}^{r}:B^{r}\to\mathbb{R}^{d\times d}$ by
    $\widetilde{g}^{r}(b)=\sum_{j=1}^{n}\rho^{r}_{j}(b)\widetilde{\Psi}^{r}_{j}(b)\widetilde{\Psi}^{r}_{j}(b)^{T}$ (32)
    and let $g^{r}=\Pi\circ\tilde{g}^{r}:B^{r}\to Gr(2,d)$.
@@
    Step 4: (Cocycle Projection) For each $j$, define $\Psi^{r}_{j}:U^{r}_{j}\to V(2,d)$ by $\Psi^{r}_{j}(b)=\widehat{\Pi}(g^{r}(b),\widetilde{\Psi}^{r}_{j}(b))$, where $\widehat{\Pi}$ is the map from Construction 3.26. Then, define $\Omega^{r}\in\check{Z}^{1}(\mathcal{U}^{r};\mathcal{C}_{O(2)})$ by $\Omega^{r}_{jk}(b)=(\Psi^{r}_{j}(b))^{T}\Psi^{r}_{k}(b)$.
@@
    $F_{j}^{r}(x)=\Psi^{r}_{j}(\pi(x))\exp\left(\sum_{k=0}^{n}\rho_{k}^{r}(\pi(x))\log\Omega_{jk}^{r}(\pi(x))f_{k}(x)\right)$ (33)
@@ 6.1 Coordinatization Pipeline For Trivial Bundles
In the special case when a family $\{f_{j}\}_{j\in J}\in T_{\alpha}(\pi,\mathcal{U}^{r})$ is witnessed by a discrete approximate coboundary $\Omega$, we construct a global trivialization map $\varphi:X\to B\times\mathbb{S}^{1}$. In particular, suppose $\omega=\text{det}_{*}(\Omega)\in\check{Z}^{1}(\mathcal{U}^{r};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}=\Delta_{\omega}(\Omega)\in\check{Z}^{2}(\mathcal{U}^{r};\mathbb{Z}_{\omega})$ are coboundaries.
@@
    Construct a solution to the linear system $\delta^{0}\phi=\omega$ over $\mathbb{Z}_{2}$. Define $\hat{\Omega}=\phi\cdot\Omega\in\check{Z}^{1}\left(\mathcal{U}^{r};\underline{SO(2)}\right)$, and let $\hat{f}_{j}=\phi_{j}\circ f_{j}$ for all $j$.
@@
Since trivial bundles are flat, we can often use Singer’s method [33] in practice to obtain a locally-constant true coboundary $\widetilde{\Omega}$ which is close to an approximate one. Though we have no theoretical guarantees on stability, the fact that the result is locally-constant guarantees that the resulting trivialization map does not have unnecessary variation.
7 Experiments
Here we demonstrate the practical viability and utility of the algorithms we have described in this paper through a series of implementations on real and synthetic datasets. A summary of the results of each experiment can be found in Table 2. Note that the roughness value $\alpha$ in each case was well outside the range required to meet our theoretical hypotheses, which suggests that our approach is stable for a larger class of objects than those covered by the present analysis. Further details about our implementation on optical flow data can be found in the companion paper [32]. Additional examples, implementation details and documentation are provided in the accompanying open-source software package.
@@ 7 Experiments
Folded Klein Bottle $\mathbb{S}^{1}$ 8 10,000 $\neq 0$ 0
Prism Densities $\mathbb{RP}^{2}$ 15,408 10,000 $\neq 0$ $\pm 3$
Table 1: A summary of the setups for the experiments. Each entry in the SW1 column indicates whether the Stiefel-Whitney class of the corresponding circle bundle model vanishes. Entries in the $\widetilde{\eta}$ column indicate the twisted Euler numbers of the models.
Experiment $\varepsilon$ $\delta$ $\alpha$ $\varepsilon_{\text{cocycle}}$ $d_{Z}(\Omega,\Pi_{Z}(\Omega))$
Optical Flow Patches 0.207 0 0.207 0 0
Folded Klein Bottle 0.207 0 0.207 0 0
Prism Densities 0.380 0.545 0.837 0.312 0.295
Table 2: Summary of experimental results. The values of $\varepsilon$, $\delta$ and $\alpha$ for each system of discrete approximate angle functions are shown. For each experiment, the constructed witness $\Omega$ was found to lie in $\check{Z}^{1}_{\varepsilon_{\text{cocycle}}}(\mathcal{U};\underline{O(2)})$.
@@
Optical Flow Patches 12 0.088 $\infty$ $\infty$ $\infty$ $\infty$ 12/12 0/0
Folded Klein Bottle 12 0.088 $\infty$ $\infty$ $\infty$ $\infty$ 12/12 0/0
Prism Densities 60 0.176 $\infty$ 0.176 $\infty$ 0.088 67/177 11/118
Table 3: Summary of persistence information for the computed characteristic class representatives in each experiment. Here $w_{\text{max}}$ denotes the maximum weight among all simplices in $\mathcal{N}(\mathcal{U})$. The cobirth and codeath weights of the Stiefel-Whitney and twisted Euler classes are respectively denoted by $b_{\omega}$, $d_{\omega}$, $b_{\widetilde{e}}$ and $d_{\widetilde{e}}$. $W^{r}$ denotes the stage of the weights filtration corresponding to $\min\{d_{\omega},d_{\widetilde{e}}\}$; the number of 1 and 2-simplices in $W^{r}$ are denoted by $(W^{r})^{(1)}$ and $(W^{r})^{(2)}$, respectively.
@@ 7.1 Optical Flow Patches
Optical flow refers to the apparent motion of objects between consecutive frames of a video [11]. This perceived motion is characterized by an assignment of a vector $v\in\mathbb{R}^{2}$ to each pixel in each frame. The vector at pixel $(x,y)$ in frame $t$, points to the location in frame $t+1$ where pixel $(x,y)$ appears to move. The estimation of optical flow from a video is central to many computer vision tasks, including motion tracking, object segmentation, and video compression [16], [34]. The Sintel dataset [4], derived from the open-source animated short film “Sintel” by the Blender Foundation, was created as a high-quality benchmark for testing and training optical flow models.
@@
In [1], Adams et al. studied the nonlinear statistics of 3 x 3 high-contrast optical flow patches from [4]. They proposed that a dense core subset $X$ of (normalized) high-contrast patches is concentrated around a 2-manifold in $\mathbb{R}^{18}$ homeomorphic to the 2-dimensional torus $\mathbb{T}^{2}$. To support their claim, they introduced a feature map $\pi_{\text{flow}}:X\to\mathbb{RP}^{1}$ which roughly measures a patch’s predominant axis of flow (see Figure 5). They then offered evidence that the data in $X$ over any small interval in $\mathbb{RP}^{1}$ has circular topology, supporting the hypothesis that an underlying circle bundle structure is present.
@@
By Proposition 2.14, there are exactly two non-isomorphic circle bundles over $\mathbb{RP}^{1}\cong\mathbb{S}^{1}$ which can be distinguished computationally by their Stiefel-Whitney classes in $H^{1}(\mathbb{RP}^{1};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$. The total spaces of the two bundles are the torus and the Klein bottle. To confirm and provide coordinates for the torus model proposed by [1], we used the same pre-processing to obtain a sample of 50,000 normalized high-contrast optical flow patches from the Sintel dataset and refine to a sample $X_{\text{flow}}$ of 25,000 patches from the dense core subset $X(300,50)$. We then computed the predominant axis of flow $\pi_{\text{flow}}(x)\in\mathbb{RP}^{1}$ for each patch $x\in X_{\text{flow}}$. Figure 6 shows the persistence diagram for a random subsample of 1,000 points from $X_{\text{flow}}$. For comparison, we also computed persistence diagrams for a random 1,000-point sampling from $X(300,30)$ and a noisy synthetic sampling of the theoretical proposed torus model. Observe that the synthetic dataset exhibits the Betti signature of $\mathbb{T}^{2}$ ($\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$) over a large range of scales, while the most persistent Betti signature for the samples of $X(300,50)$ and $X(300,30)$ is $\beta_{0}=1$, $\beta_{1}=1$, $\beta_{2}=0$.
Figure 6: Persistence diagrams for the top 50% and 30% densest subsets of the original dataset (left, middle), and those of a synthetically generated dataset from the hypothesized torus model (right).
On the other hand, as observed in [1], the local picture of $X_{\text{flow}}$ with respect to $\pi_{\text{flow}}$ tells a different story. We covered $\mathbb{RP}^{1}$ with 12 open intervals of the form $U_{j}=(\frac{j\pi}{12}-\frac{\pi}{16},\frac{j\pi}{12}+\frac{\pi}{16})$ (note that $\frac{\pi}{16}$ is the maximal radius for which the nerve of $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ has no triple intersections). Figure 7 shows the first two principal components and the persistence diagrams for the data in each $\pi_{\text{flow}}^{-1}(U_{j})$, strongly suggesting the presence of an underlying circle bundle structure.
@@
We computed a family $\{f_{j}\}_{j=1}^{12}$ of discrete approximate local trivializations for $X_{\text{flow}}$ subordinate to $\mathcal{U}$ using PCA, then obtained a witness $\Omega\in\check{C}^{1}(\mathcal{N}(\mathcal{U});O(2))$ using the procedure outlined in Construction 3.16. Figure 8 shows a visualization of $\mathcal{N}(\mathcal{U})$; the weight $w(jk)$ of each edge $(jk)$ is shown in black, and the values assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)$ are shown in red. A 0-cochain $\phi\in C^{0}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ satisfying $\delta^{0}\phi=\omega$ is shown in blue, confirming a topological torus structure underlying the data.
Figure 8: (Left) Correlations on intersections between the discrete approximate local trivializations $\{f_{j}\}_{j=1}^{12}$ computed for $X_{\text{flow}}$. (Right) The nerve of the open cover $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ of $\mathbb{RP}^{1}$. Each edge is labeled by its associated weight. The components of the Stiefel-Whitney class representative $\omega$ are shown in blue, and a potential $\phi$ for $\omega$ is shown in red.
Once we verified the global triviality of the underlying bundle structure, we used the coordinatization pipeline described in Section 6 to obtain a global coordinate system for $X_{\text{flow}}$. It is important to note that a significant portion of the patches had ambiguous predominant flow directions despite being high-contrast (see Figure 9). Such patches are not accounted for in the torus model proposed by [1]. Upon further investigation, we found that the fibers of the predominant direction map $\pi_{\text{flow}}$ are geometrically more like cylinders than circles, and that this extra degree of freedom locally parametrizes variations in directional strength among patches in $X_{\text{flow}}$. The generalized model we describe in our upcoming work also accounts for other families of high-contrast patches not featured in the original torus model which are particularly relevant for motion tracking analysis.
@@
Figure 10: A sample of coordinatized patches from $X_{\text{flow}}$ with high directional strength, arranged by predominant flow direction (x) and assigned fiber coordinate (y). Compare with Figure 7 in [1]
From a modeling perspective, the results of our analysis of $X_{\text{flow}}$ demonstrate how the interplay between local and global measurements depends crucially on the geometry of the chosen feature map. On the other hand, a careful fiberwise analysis helped us to both confirm an existing model and identify subtle, informative additional structure in the data which had previously been missed. For additional details, see our companion paper [32].
7.2 Folded Klein Bottle
We demonstrate our pipeline on a noisy synthetic sampling of a 2-manifold $M\subset\mathbb{R}^{8}$ with the topology of a Klein bottle. Unlike our previous example, the dataset does not come equipped with any obvious choice of projection map, so we use the DREiMac library’s topologically-flavored circular coordinates algorithm [19], [20] to obtain an $\mathbb{S}^{1}$-valued base map. The data in the fibers of the resulting projection map is concentrated around topological (but not geometric) circles. We therefore also use the DREiMac algorithm to construct local circular coordinates.
The model manifold $M$ is defined as follows: let $\gamma:\mathbb{R}\to\mathbb{R}^{4}$ be defined by $\gamma(t)=(\sin(t),\cos(t),\sin(2t),0)$ for all $t\in\mathbb{R}$, and let $F=\gamma(\mathbb{R})\subset\mathbb{R}^{4}$. Note that $F$ is topologically a circle, but geometrically it is folded. In particular, $F$ is invariant under a reflection in the $x$ and $y$ coordinates. Now, define $R:[0,2\pi]\to SO(4)$ by
$R(t)=\left(\begin{array}[]{cccc}a&b&0&d\\ b&a&0&-d\\ 0&0&1&0\\ -d&d&0&c\\ \end{array}\right)$
where
$\tau=\frac{t}{2},\hskip 28.45274pta=\frac{1+\cos(\tau)}{2},\hskip 28.45274ptb=\frac{1-\cos(\tau)}{2}$
$c=\cos(\tau),\hskip 28.45274ptd=\frac{\sqrt{2}}{2}\sin(\tau)$
Observe in particular that $R(0)=I$, $R(2\pi)$ is the matrix which maps $(x,y,z,w)$ to (y,x,z,w)(y,x,z,-w) and $R(t)$ fixes the $z$-axis for all $t$. Finally, define $\widetilde{M}$ by
$\widetilde{M}=\{(u,v)\in\mathbb{R}^{4}\times\mathbb{R}^{4}:u=\gamma(t),\ v\in R(t)F\}$
and apply an $SO(8)$ rotation to obtain the manifold $M$ (the rotation is simply to intermix the coordinates).
We generated a dataset $X_{\text{kb}}\subset\mathbb{R}^{8}$ by sampling $10,000$ points from $M$ with a Gaussian noise level of $\sigma=0.05$. Figure 11 shows persistence diagrams with $\mathbb{Z}_{2}$- and $\mathbb{Z}_{3}$-coefficients for a subsample of $X_{\text{kb}}$ (recall that the Klein bottle has Betti numbers $\beta_{0}=0$, $\beta_{1}=1$, $\beta_{2}=0$, with 2-torsion classes in homology dimensions 1 and 2):
Figure 11: Persistence diagrams for a subsample of $X_{kb}$
We used the $\mathbb{Z}_{3}$-persistent class as input for the DREiMac circular coordinates algorithm to produce a map $\pi_{\text{kb}}:X_{\text{kb}}\to\mathbb{S}^{1}$. We then constructed a cover of $\mathbb{S}^{1}$ by metric balls $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ around equally-spaced landmarks. Figure 12 shows PCA projections of the data in $\pi^{-1}_{\text{kb}}(U_{j})$ for several $j$, confirming that PCA could not be used to produce local circular coordinates. The points are colored according to angular coordinates computed using DREiMac:
Figure 12: PCA projections of $\pi_{\text{kb}}^{-1}(U_{j})$ for several $j$.
A characteristic class computation then confirmed the anticipated global Klein bottle structure; a visualization of the nerve decorated by the non-trivial Stiefel-Whitney class representative is shown in Figure 13.
Figure 13: The nerve of the open cover $\mathcal{U}=\{U_{j}\}_{j=1}^{12}$ of $\mathbb{S}^{1}$. Each edge is labeled by its associated weight. The components of the Stiefel-Whitney class representative $\omega$ are shown in blue. Note the class is non-trivial on the full nerve, reflecting the underlying Klein bottle structure. After the heaviest edge is removed, the class becomes trivial; a potential for the restricted cocycle is shown in red.